# QAPyramid: Fine-grained Evaluation of Content Selection for Text Summarization

### v1

@@ 3.4 Result Analysis
Table 1 shows the metric scores of 10 summarization systems on the 50-example subset from the CNN/DM test set (and Table 6 in the Appendix contains systems scores of many other automatic or semi-automatic metrics). The ACU and $n$ACU scores are computed based on the raw annotation data released by Liu et al. (2023b). [^3] https://huggingface.co/datasets/Salesforce/rose
First, a consistent trend across different metrics is that BRIO and BRIO-Ext are the two best-performing models, likely because they are carefully fine-tuned on the CNN/DM training set.
Second, 1-shot LLMs consistently obtain worse ROUGE-2 recall or F1 scores than fine-tuned models. However, their QAPyramid and $n$QAPyramid scores are comparable to those of fine-tuned models. This demonstrates that, compared to metrics based on lexical overlaps, our QAPyramid method captures more semantic similarities and thus alleviates the rigidity caused by reference-based evaluations with a single reference. This finding is also consistent with previous observations that zero-shot or few-shot LLMs are preferred by humans despite their low ROUGE scores Goyal et al. (2023).
Third, we observe that QAPyramid scores are overall higher than ACU scores. One hypothesis is that QAs are finer-grained than ACUs, so they give credit to finer-grained alignments between the system and reference summaries. To support this hypothesis, we find that in 21% of cases, a predicate only has a subset of its QA pairs present in the system summary, i.e., if the judgment is at the ACU level, it would miss some partial correctness. Next, we randomly sampled examples from summaries generated by fine-tuned models and manually compared QAPyramid with ACU. We find that, besides QA pairs being finer-grained, two other factors also contribute to higher QAPyramid scores: (1) our annotation guideline led to more lenient judgments of “being present” based more on semantics than lexicon, and (2) the predicate-centered nature of QA-SRL may cause one piece of information to be credited multiple times. See Appendix B for two examples.
@@ 4 QAPyramid Automation
Here we explore how to automate QAPyramid, where we break it down into automation of QA generation and QA presence detection.
QA Generation
Method RL BS UA
QASem (flan-t5-large) 78.9 94.5 99.9
QASem (flan-t5-xl) 79.0 94.5 99.9
Llama-3.1-8B-Inst 0shot 36.9 87.9 24.6
1shot 49.4 90.0 47.7
3shot 56.9 91.2 57.2
5shot 61.2 91.9 60.8
Llama-3.1-70B-Inst 0shot 41.3 87.1 32.1
1shot 61.3 91.3 66.1
3shot 69.4 92.8 72.7
5shot 72.0 93.3 74.6
GPT-4o 0shot 45.9 88.9 41.5
1shot 62.0 91.4 71.0
3shot 69.4 92.7 76.7
5shot 72.2 93.3 78.2
Table 2: Automatic QA generation performance. RL, BS, and UA refer to ROUGE-L, BERTScore, and unlabeled argument detection, respectively.
@@ 4.1 QA Generation Automation
We explore two types of models to conduct this task automatically. First, we consider prompting large language models (LLMs). We use open-sourced LLMs, Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct Dubey et al. (2024), as well as a proprietary model, GPT-4o OpenAI (2024). We test with 0, 1, 3, and 5 in-context examples [^4] In our preliminary experiments, we do not find any significant gains using more than 5 in-context examples. randomly sampled from the set. To prevent answer leakage, we ensure that the in-context examples do not contain the same reference summary as the test example. LLM prompts are shown in Figure 7. Second, we use a fine-tuned model, QASem Klein et al. (2022), which was jointly trained on the QASRL FitzGerald et al. (2018) and QANom Klein et al. (2020) datasets based on T5 models. The parser takes as input a sentence and a predicate and generates a set of QA pairs. In our experiments, we use an improved version of the parser Cattan et al. (2024), which leverages Flan-T5-Large and Flan-T5-XL Shen et al. (2023). [^5] https://github.com/plroit/qasem_parser
To evaluate how similar the generated QAs are to those written by humans, we assess the similarity between the generated and gold QA pairs using ROUGE-L Lin (2004), BERTScore Zhang* et al. (2020), and a previously developed metric from QASRL Roit et al. (2020) and QANom Klein et al. (2020) tasks, Unlabeled Argument Detection (UA). UA calculates token overlap between the generated and gold answers. [^6] We exclude labeled argument detection, the metric on question equivalence because it relies on a rule-based parser that works only on questions that follow predefined automata and is therefore not suitable for our “free-form” questions, as mentioned in Section 3.1.
We report the results in Table 2. For the LLM-based models, we observe a clear trend: having more in-context examples consistently improves performance. For example, the UA metric is more than doubled from zero-shot to five-shot settings. Within the Llama-3.1 models, the larger model (70B) outperforms the smaller model (8B) on all metrics and in all in-context settings, except for the zero-shot setting with BERTScore. Interestingly, the performance of Llama-3.1-70B-Instruct is quite comparable to that of GPT-4o, achieving similar BERTScore and ROUGE-L scores. This demonstrates the promise of using open-source models for this QA generation task. Finally, we observe that the two fine-tuned models achieve the highest scores, which is expected since they have been trained to generate QAs of similar styles. Thus, we recommend using the fine-tuned QASem model to automate the QA generation step.
Method QA F1 Stmt F1
DeBERTa v3 81.6 79.3
MiniCheck 78.8 80.1
@@
Llama-3.1-8B-Inst 0shot 53.8 69.6
1shot 79.9 77.4
3shot 81.1 78.5
5shot 81.5 78.3
Llama-3.1-70B-Inst 0shot 81.5 80.4
1shot 84.9 80.6
3shot 84.8 82.3
5shot 84.7 83.0
GPT-4o 0shot 78.8 81.8
1shot 82.9 83.5
3shot 84.7 83.7
5shot 85.0 84.2
Table 3: Automatic QA presence detection performance. We report micro F1 scores in two scenarios where we take the QA pair as is or transform it into a statement (Stmt).
@@
4.2 QA Presence Detection Automation
Similar to QA generation, we explore LLM-based models with different numbers of in-context examples, excluding examples that use the same reference summary. Since detecting the presence of QAs is a typical natural language inference (NLI) task, we explore an off-the-shelf pretrained NLI model, DeBERTa v3888We use the model that has been fine-tuned on many tasks for strong classification performance: sileod/deberta-v3-base-tasksource-nli. He et al. (2023). Additionally, we test two NLI-based faithfulness evaluation metrics for summarization: MiniCheck Tang et al. (2024) and AlignScore Zha et al. (2023). They were found to highly correlate with human judgments in whether a system summary is entailed by the document. In typical entailment tasks, the hypothesis is usually a statement rather than a QA. Therefore, we also explore transforming the QAs into statements using GPT-4o. Finally, we report the micro F1 score to assess how well the predicted labels match the gold labels.
@@ 4.3 Meta Evaluation of Automated Metrics
Since we believe QAPyramid provides more reliable and accurate human evaluation signals, we use it to benchmark automated metrics. We test the correlation between QAPyramid and another Pyramid-style human evaluation protocol, ACU Liu et al. (2023b), and its automated version, AutoACU Liu et al. (2023c) (we use A2CU). Although Liu et al. (2023c) did not explore the semi-automatic option, we test SemiAutoACU metric using their human-written atomic units and their trained unit presence checking model. We also include metrics introduced by Zhang and Bansal (2021) that semi-automate or fully automate the Pyramid method, namely Lite2Pyramid and Lite3Pyramid, respectively. For Lite2Pyramid, we utilize the gold units from ACU (because SCUs are not available for the 50 examples in our dataset) while employing Zhang and Bansal (2021) trained presence detection model.
@@
Lastly, we include various widely adopted summarization evaluation metrics: ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), CHRF Popović (2015), BERTScore Zhang* et al. (2020), BARTScore Yuan et al. (2021), and a variant of GEval Liu et al. (2023a) [^9] We adapted the prompt from Liu et al. (2023b) for this task., a GPT-4-based metric that predicts a numerical score. Since QAPyramid is recall-focused, we use the recall version of these metrics when available.
@@
Following previous works Fabbri et al. (2021); Liu et al. (2023b), we use system-level and summary-level Kendall’s tau correlations as our meta-evaluation metrics. The results are presented in Table 4. We report the results separately and collectively for the 5 fine-tuned (FT) models and the 5 LLM-based models. First, for both ACU and QAPyramid, their semi-automatic metrics outperform their fully automatic counterparts. Surprisingly, Lite3Pyramid works better than Lite2Pyramid with ACU perhaps due to the mismatch between ACUs and the original SCUs used by Lite2Pyramid. Second, we find that AutoACU or Lite3Pyramid, which uses ACUs or SCUs, has lower correlations than AutoQAPyramid. This is due to the mismatch between their original Pyramid-style human evaluations and QAPyramid. Nonetheless, we believe QAPyramid is a more reliable protocol and it is desired to best correlate with it. Lite3Pyramid correlates surprisingly well with QAPyramid probably because their automated SCUs (called STUs) are extracted based on SRL. Overall, our SemiAutoPyramid and AutoQAPyramid metrics show higher correlations with QAPyramid than other metrics. This demonstrates that they effectively automate QAPyramid, thus providing finer-grained automatic evaluation methods for text summarization.
5 Related Works
Human Evaluation for Text Summarization.
Human evaluation is the ultimate way to assess the quality of model-generated summaries. However, there is no consensus on how human evaluation should be conducted. In many cases, people resort to direct rating, i.e., humans directly rate the quality of the summary (or rate for certain aspects, e.g., relevance Fabbri et al. (2021)). However, direct rating often suffers from subjectivity, low agreement, and thus non-reproducibility Kiritchenko and Mohammad (2017); Falke et al. (2019); Shapira et al. (2019). Two factors contribute to this issue: (1) the summary is usually more than one sentence and different annotations may put their focus on different parts of the summary; (2) when reference is not provided, what content is considered important and should be selected into the summary may vary from person to person. The canonical Pyramid Nenkova and Passonneau (2004) method was introduced to improve reproducibility. First, it is reference-based, which resolves the question of what content needs to be selected. Second, it decomposes the summary into smaller units, which reduces the ambiguity when evaluating long text. The original Pyramid method requires multiple references and full expert annotations. Shapira et al. (2019) simplified Pyramid, making it feasible for single reference, and made the judgment of unit presence crowdsourceable. Liu et al. (2023b) introduced the ACU (Atomic Content Unit) protocol that further clarified the concept of atomic units. However, there are still no standard guidelines for ACU or SCU. How to write each unit and what content to include in each unit are purely dependent on the annotator. Getting different sets of units from different annotators compromises the reproducibility of this protocol. In contrast, the units (QAs) of QAPyramid are systematically defined by QA-SRL. Furthermore, QA-SRL-based QAs are usually finer-grained than ACUs or SCUs. Another human evaluation approachis best-worst scaling or pairwise comparison Kiritchenko and Mohammad (2017), which has become a standard way to benchmark modern LLMs Goyal et al. (2023); Chiang et al. (2024). However, it also suffers from the low-agreement issue when text is long and it scales quadratically with the number of systems.
Automatic Evaluation for Text Summarization.
Compared to human evaluations, automatic metrics are fast, cheap, and reproducible. However, how well they correlate with human judgment is always a concern. Over the years, many automatic evaluation metrics have been introduced. Some early metrics measure the n-gram overlap between system and reference summaries Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005), out of which ROUGE Lin (2004) is the most widely adopted metric till today. To alleviate the rigidity of exact lexical match, metrics based on the similarity between embeddings were introduced Zhao et al. (2019); Clark et al. (2019); Zhang* et al. (2020). There are also works automating the Pyramid protocol Yang et al. (2016); Hirao et al. (2018); Gao et al. (2019); Zhang and Bansal (2021); Liu et al. (2023c); Nawrath et al. (2024). They use Open IE, SRL, AMR, fine-tuned models, or LLMs to automate unit extraction and use NLI models to automate unit presence detection. Compared to these works, AutoQAPyramid is advantageous because it automates a more systermatical, reproducible, and fine-grained human evaluation protocol, QAPyramid. Besides evaluating content selection, a separate line of research has been focused on faithfulness or factuality evaluation, i.e., checking if the source document(s) entail the summary Cao et al. (2018); Maynez et al. (2020); Laban et al. (2022); Zha et al. (2023). The Pyramid type of methods have also been applied in this scenario Chen et al. (2023); Min et al. (2023); Kamoi et al. (2023); Wan et al. (2024); Tang et al. (2024); Gunjal and Durrett (2024) and the units they extract are ACU or SCU style units. There are some methods based on question generation and answering Wang et al. (2020); Durmus et al. (2020); Scialom et al. (2021); Fabbri et al. (2022). However, their QAs are not QA-SRL-based questions and each QA contains more than one argument. Except for a most recent work Cattan et al. (2024), it adopted QA-SRL to decompose system summary automatically and manually verify each QA against the source document. Technically, our AutoQAPyramid methods can also be applied for faithfulness evaluation by generating QAs from the system summary and checking QA presence against the source document.
@@ 6 Conclusion
Limitation
One limitation of QAPyramid is that we considered only QA-SRL He et al. (2015); FitzGerald et al. (2018) when generating QA pairs. As mentioned in Section 2.2, QA pairs can also be generated for deverbal nominalizations Klein et al. (2020), discourse relations Pyatkin et al. (2020), and adjectival semantics Pesahov et al. (2023). Future work can focus on how to more thoroughly decompose the meaning of the reference summary.
Another limitation of QAPyramid is that it is only applicable to reference-based evaluation, a constraint shared by all Pyramid-style evaluations. Such evaluation requires human-written references, which may not always be available. Moreover, when only a single reference is provided, it must accurately capture all the content that should be included. If multiple different ways of content selection are all valid, an evaluation based on a single reference may be biased. Despite this rigidity, reference-based evaluation is usually more reproducible and controllable compared to reference-free evaluation Deutsch et al. (2022).
@@ Acknowledgments
We thank Paul Roit, Ori Shapira, and Ori Ernst for the helpful discussions. This work was supported by NSF-CAREER Award 1846185, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, Microsoft Accelerate Foundation Models Research (AFMR) grant program, a Google PhD Fellowship, the Israel Science Foundation (grant no. 2827/21), by a grant from the Israeli Planning and Budgeting Committee (PBC), and the PBC fellowship for outstanding PhD candidates in data science. The views contained in this article are those of the authors and not of the funding agency.

### v2

@@ 3.4 Result Analysis
Table 1 shows the metric scores of 10 summarization systems on the 50-example subset from the CNN/DM test set (and Table 6 in the Appendix contains systems scores of many other automatic or semi-automatic metrics). The ACU and nnACU scores are computed based on the raw annotation data released by Liu et al. (2023b). [^4] https://huggingface.co/datasets/Salesforce/rose First, a consistent trend across different metrics is that BRIO and BRIO-Ext are the two best-performing models, likely because they are carefully fine-tuned on the CNN/DM training set. Second, 1-shot LLMs consistently obtain worse ROUGE-2 recall or F1 scores than fine-tuned models. However, their QAPyramid and nnQAPyramid scores are comparable to those of fine-tuned models. This demonstrates that, compared to metrics based on lexical overlaps, our QAPyramid method captures more semantic similarities and thus alleviates the rigidity caused by reference-based evaluations with a single reference. This finding is also consistent with previous observations that zero-shot or few-shot LLMs are preferred by humans despite their low ROUGE scores (Goyal et al., 2023). Third, we observe that QAPyramid scores are overall higher than ACU scores. One hypothesis is that QAs are finer-grained than ACUs, so they give credit to finer-grained alignments between the system and reference summaries. To support this hypothesis, we find that in 21% of cases, a predicate only has a subset of its QA pairs present in the system summary, i.e., if the judgment is at the ACU level, it would miss some partial correctness. Next, we manually examined some examples and we find that, besides QA pairs being finer-grained, two other factors also contribute to higher QAPyramid scores: (1) our annotation guideline led to more lenient judgments of “being present” based more on semantics than lexicon, and (2) the predicate-centered nature of QA-SRL may cause one piece of information to be credited multiple times. See Appendix C for two examples.
@@ 4.1 QA Generation Automation
We explore two types of models to conduct this task automatically. First, we consider prompting large language models (LLMs). We use open-sourced LLMs, Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct (Llama Team et al., 2024), as well as a proprietary model, GPT-4o (OpenAI, 2024). We test with 0, 1, 3, and 5 in-context examples [^5] We do not find any significant gains using more than 5 in-context examples. randomly sampled from the set. To prevent answer leakage, we ensure that the in-context examples do not contain the same reference summary as the test example. LLM prompts are shown in Figure 7. Second, we use a fine-tuned model, QASem (Klein et al., 2022), which was jointly trained on the QASRL (FitzGerald et al., 2018) and QANom (Klein et al., 2020) datasets based on T5 models. The parser takes as input a sentence and a predicate and generates a set of QA pairs. In our experiments, we use an improved version of the parser (Cattan et al., 2024), which leverages Flan-T5-Large and Flan-T5-XL (Shen et al., 2023). [^6] https://github.com/plroit/qasem_parser
To evaluate how similar the generated QAs are to those written by humans, we assess the similarity between the generated and gold QA pairs using ROUGE-L (Lin, 2004), BERTScore (Zhang* et al., 2020), and a previously developed metric from QASRL (Roit et al., 2020) and QANom (Klein et al., 2020) tasks, Unlabeled Argument Detection (UA). UA calculates token overlap between the generated and gold answers. [^7] We exclude labeled argument detection, the metric on question equivalence because it relies on a rule-based parser that works only on questions that follow predefined automata and is therefore not suitable for our “free-form” questions, as mentioned in Section 3.1.
We report the results in Table 2. For the LLM-based models, we observe a clear trend: having more in-context examples consistently improves performance. For example, the UA metric is more than doubled from zero-shot to five-shot settings. Within the Llama-3.1 models, the larger model (70B) outperforms the smaller model (8B) on all metrics and in all in-context settings, except for the zero-shot setting with BERTScore. Interestingly, the performance of Llama-3.1-70B-Instruct is quite comparable to that of GPT-4o, achieving similar BERTScore and ROUGE-L scores. This demonstrates the promise of using open-source models for this QA generation task. Finally, we observe that the two fine-tuned models achieve the highest scores, which is expected since they have been trained to generate QAs of similar styles. Thus, we recommend using fine-tuned QASem models to automate the QA generation step.
QA Generation QA Presence
Method RL BS UA Method QA F1 Stmt F1
QASem (flan-t5-large) 78.9 94.5 99.9 DeBERTa v3 81.6 79.3
QASem (flan-t5-xl) 79.0 94.5 99.9 MiniCheck 78.8 80.1
@@
Llama-3.1-8B-Inst 0shot 36.9 87.9 24.6 Llama-3.1-8B-Inst 0shot 53.8 69.6
1shot 49.4 90.0 47.7 1shot 79.9 77.4
3shot 56.9 91.2 57.2 3shot 84.8 82.3
5shot 61.2 91.9 60.8 5shot 81.5 78.3
Llama-3.1-70B-Inst 0shot 41.3 87.1 32.1 Llama-3.1-70B-Inst 0shot 81.5 80.4
1shot 61.3 91.3 66.1 1shot 84.9 80.6
3shot 69.4 92.8 72.7 3shot 84.8 82.3
5shot 72.0 93.3 74.6 5shot 84.7 83.0
GPT-4o 0shot 45.9 88.9 41.5 GPT-4o 0shot 78.8 81.8
1shot 62.0 91.4 71.0 1shot 82.9 83.5
3shot 69.4 92.7 76.7 3shot 84.7 83.7
5shot 72.2 93.3 78.2 5shot 85.0 84.2
Table 2: Automatic QA generation and presence detection performance. RL, BS, and UA refer to ROUGE-L, BERTScore, and unlabeled argument detection, respectively. For QA presence detection, we report micro F1 scores in two scenarios where we take the QA pair as is or transform it into a statement (Stmt).
4.2 QA Presence Detection Automation
Similar to QA generation, we explore LLM-based models with different numbers of in-context examples, excluding examples that use the same reference summary. Since detecting the presence of QAs is a typical natural language inference (NLI) task, we explore an off-the-shelf pretrained NLI model, DeBERTa v3999We use the model that has been fine-tuned on many tasks for strong classification performance: sileod/deberta-v3-base-tasksource-nli. (He et al., 2023). Additionally, we test two NLI-based faithfulness evaluation metrics for summarization: MiniCheck (Tang et al., 2024) and AlignScore (Zha et al., 2023). They were found to highly correlate with human judgments in whether a system summary is entailed by the document. In typical entailment tasks, the hypothesis is usually a statement rather than a QA. Therefore, we also explore transforming the QAs into statements using GPT-4o. Finally, we report the micro F1 score to assess how well the predicted labels match the gold labels.
@@ 4.3 Meta Evaluation of Automated Metrics
Since we believe QAPyramid provides more reliable and accurate human evaluation signals, we use it to benchmark automated metrics. We test the correlation between QAPyramid and another Pyramid-style human evaluation protocol, ACU (Liu et al., 2023b), and its automated version, AutoACU (Liu et al., 2023c) (we use A2CU). Although Liu et al. (2023c) did not explore the semi-automatic option, we test SemiAutoACU metric using their human-written atomic units and their trained unit presence checking model. We also include metrics introduced by Zhang & Bansal (2021) that semi-automate or fully automate the Pyramid method, namely Lite2Pyramid and Lite3Pyramid, respectively. For Lite2Pyramid, we utilize the gold units from ACU (because SCUs are not available for the 50 examples in our dataset) while employing Zhang & Bansal (2021) trained presence detection model.
@@
Lastly, we include various widely adopted summarization evaluation metrics: ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), CHRF (Popović, 2015), BERTScore (Zhang* et al., 2020), BARTScore (Yuan et al., 2021), and a variant of GEval (Liu et al., 2023a) [^10] We adapted the prompt from Liu et al. (2023b) for this task., a GPT-4-based metric that predicts a numerical score. Since QAPyramid is recall-focused, we use the recall version of these metrics when available.
@@
Following previous works (Fabbri et al., 2021b; Liu et al., 2023b), we use system-level and summary-level Kendall’s tau correlations as our meta-evaluation metrics. The results are presented in Table 3. We report the results separately and collectively for the 5 fine-tuned (FT) models and the 5 LLM-based models. First, for both ACU and QAPyramid, their semi-automatic metrics outperform their fully automatic counterparts. Surprisingly, Lite3Pyramid works better than Lite2Pyramid with ACU perhaps due to the mismatch between ACUs and the original SCUs used by Lite2Pyramid. Second, we find that AutoACU or Lite3Pyramid, which uses ACUs or SCUs, has lower correlations than AutoQAPyramid. This is due to the mismatch between their original Pyramid-style human evaluations and QAPyramid. Nonetheless, we believe QAPyramid is a more reliable protocol and it is desired to best correlate with it. Lite3Pyramid correlates surprisingly well with QAPyramid probably because their automated SCUs (called STUs) are extracted based on SRL. Overall, our SemiAutoPyramid and AutoQAPyramid metrics show higher correlations with QAPyramid than other metrics. This demonstrates that they effectively automate QAPyramid, thus providing finer-grained automatic evaluation methods for text summarization.
@@ 5 Conclusion
6 Limitations
First, our evaluation is confined to the English CNN/DM news summarization dataset; the generalizability of our findings to other languages and domains remains to be explored.
Second, as a Pyramid-style metric, QAPyramid is strictly reference-based. This reliance on human-authored references can be a constraint, as they may be unavailable or biased if only a single reference is provided. This limitation is evident in our experiments on the reference-free SummEval benchmark (Appendix A.1), where we observe weak correlations. The discrepancy arises because our metric is tightly grounded in the reference summary, whereas human ratings often permit greater flexibility in content selection.
Finally, our semantic decomposition could be more comprehensive. We currently use only QA-SRL (He et al., 2015; FitzGerald et al., 2018) to generate question-answer (QA) pairs, but future work could incorporate methods for nominalizations (Klein et al., 2020) or discourse relations (Pyatkin et al., 2020) for a more thorough analysis. Relatedly, our metric weights all QA pairs equally, which may allow numerous minor arguments to overshadow more salient information. This could be addressed by exploring predicate-level aggregation, where scores are averaged within a predicate before being combined.
@@ Acknowledgments
We thank the anonymous reviewers for their helpful comments, and we also thank Paul Roit, Ori Shapira, and Ori Ernst for the helpful discussions. This work was supported by NSF-CAREER Award 1846185, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, Microsoft Accelerate Foundation Models Research (AFMR) grant program, a Google PhD Fellowship, the Israel Science Foundation (grant no. 2827/21), by a grant from the Israeli Planning and Budgeting Committee (PBC), and the PBC fellowship for outstanding PhD candidates in data science. The views contained in this article are those of the authors and not of the funding agency.